OpenStudy (anonymous):

Under certain conditions argon (Ar) gas diffuses at a rate of 3.2 centimeters per second. Under the same conditions, an unknown gas diffuses at a rate of 4.5 centimeters per second. What is the approximate molar mass of the unknown gas? 10 g/mol 20 g/mol 28 g/mol 56 g/mol

OpenStudy (ipwnbunnies):

\[\frac{rate_A}{rate_B} = \sqrt{\frac{M_b}{M_A}}\]

OpenStudy (ipwnbunnies):

Where 'M' is the molar mass of the substance. Let 'A' be the argon and 'B' be the unknown substance. Solve for Mb, which would be the molar mass of the unknown substance.

OpenStudy (anonymous):

@iPwnBunnies how would you get the mol. i know how to plug in the first set of numbers (the rates) but i dont know how to complete the rest.

OpenStudy (ipwnbunnies):

We're solving for Mb. Ma will be the molar mass, or atomic mass, of the Argon, which is found in the Periodic Table.

OpenStudy (ipwnbunnies):

It's 39.95 g/mol or Daltons, or w/e. The units don't matter in this problem because the units will cancel out anyway.

OpenStudy (anonymous):

@iPwnBunnies thank you so much i got it now.(: